Minimum distance. Find the minimum distance from a point on the

positive x-axis (a, 0) to the parabola y^2 = 8x.


1
Expert's answer
2021-12-20T05:23:08-0500

Given parabola y² = 8x


and the point (a,0).


Let P(x, y) be any point on the parabola.


"d = \sqrt{(x − a ) ² + y²}"


"D=d\u00b2 = (x-a)\u00b2 + y\u00b2"


"D(x)= (x \u2212 a)\u00b2 +8x"

Differentiate with respect to x


"D'(x) = 2(x-a)+8"


"D'(x) = 0"


"2(x-a)+8=0"


"x-a=-4"


"x = -4+a"

"D'(x) = 2(x-a)+8"


Differentiate with respect to x


"D''(x) = 2x"


"D''(-4+a)= 2(-4+a)= -8+a"


If a > 8 then D"(-4+a) >0


The minimum distance is

"\sqrt{(-4+a-a)² +8(-4+a)}"


="\sqrt{16-12+8a}"


= "\sqrt{4+8a}"